the straight line 2x-y+c=0 is the tangent to the ellipse 4x^2+8y^2=36 from ( (5)^0.5,0) and ( -(5)^0.5,0) is?

Grade:12th pass

1 Answers

Vikas TU
6 years ago
 The Question doesn’t ask to derive or solve for anything.
The equation of the tangent to x2/a2 + y2/b2 = 1 at P(x1,y1) is xx1/a2 + yy1/b2 = 1.
Since ( (5)^0.5,0) is point on the line
2((5)^0.5) – 0 + c =0
c = - 2root5
Hence, equation is 2x – y - 2root5 = 0.